F(x) = √(tanx) + √(sinx)|³|²
1
Expert's answer
2020-06-08T18:23:34-0400
"F(x)=\sqrt{\tan(x)}+\sqrt[3]{\sin^2(x)}"

"F'(x)={1\over 2\sqrt{\tan(x)}}(\tan(x))'+{2\over 3\sqrt[3]{\sin(x)}} (\sin(x))'="

"={1\over 2\cos^2(x)\sqrt{\tan(x)}}+{2\cos(x)\over 3\sqrt[3]{\sin(x)}}"